An electron is located on the x axis at x0 = -4.91 × 10-6 m. Find the magnitude and direction of the electric field at x = 1.43 × 10-6 m on the x axis due to this electron.

Magnitude:______N/C

Direction:

a. Positive x direction

b. Negative x direction

c. none of these
Answer #1

E = KQ/d^2

HERE d = 6.34*10^-6

E = 9*10^9*1.6*10^-19/(6.34*10^-6)^2

E = 35.8 N/C

direction field is action along the negative direction